1-A camera uses a 200.0-mm focal length telephoto lens to take pictures from a distance of infinity to as close as 2.0 m. What are the minimum and maximum distances from the lens to the sensor?
1
Expert's answer
2020-04-02T10:44:02-0400

The minimum distances


"\frac{1}{a}+\frac{1}{b}=\frac{1}{f}\to\frac{1}{\infin}+\frac{1}{b}=\frac{1}{0.2}\to b=0.2m"


The maximum distances


"\frac{1}{a}+\frac{1}{b}=\frac{1}{f}\to\frac{1}{2}+\frac{1}{b}=\frac{1}{0.2}\to b=0.222m"